# MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

### v2

@@ 3.5 Ablation Study
To evaluate the effectiveness of each component of our LoT-TL training pipeline, we conduct this thorough ablation study. We demonstrate that the elements in LoT-TL pipeline work synergistically to strengthen the model’s formal theorem-proving capability through the integration of FL in Long CoT. We use DeepSeek-Prover-v1.5-SFT as our base model and apply the pass@64 accuracy rate on the whole-proof generation method for this set of experiments. The results are presented in Table 3.
@@ 3.5.1 Effect of Training Stages
We evaluate training progression by measuring performance across key intermediate models, namely base model, base model + Long CoT, and base model + Long CoT + SFT, as shown in Table 3. Results indicate that training solely on NL CoT data provides minimal improvement, suggesting that NL CoT reasoning alone does not make Lean CoT capability emerge. However, incorporating SFT data with the LoT-TL training method yields a marked improvement, demonstrating the effectiveness of transfer learning in equipping models with Lean4 Long CoT capabilities. Interestingly, additional training with correction data, although it is not designed for whole-proof writing, further enhances the performance. This improvement likely arises from the model developing self-analysis capabilities of Lean4 code, allowing it to avoid potentially wrong solutions.
@@ 3.5.2 Switch-off Long CoT
This experiment shows that the strong FL reasoning power of LoT-Solver comes from the emergent formal reasoning ability in Long CoT, rather than trivially stacking more data. It uses our LoT-Solver model to write Lean4 proofs directly using the code completion method without Long CoT. We find the performance drop from 51.64% to 49.18%, which is because the model does not take explicit high-level plan in Long CoT, making it unable to finish some questions in the induction field.
@@ 3.5.3 Ablation of Long CoT
To validate the quality of integration between NL and FL in Long CoT, we fine-tune the DeepSeek-Prover-v1.5-RL model directly using our SFT dataset without Long CoT reasoning. We can find that the performance of w/o Long CoT model (48.36%) is lower than LoT-Solver (51.64%). This confirms that Long CoT plays a crucial role in Lean4 theorem proving, offering structured reasoning that outperforms direct RL-based additional data fine-tuned model.
@@ 3.6 Case Study
This section presents the general results of case studies of MA-LoT framework. Due to the limited space, we leave detailed examples in Appendix D. The results show the collaboration between agents. The prover agent can use a high-level plan to prove advanced MATH theorems and the corrector can analyze the feedback from the Lean executor to formulate correct proof of IMO-level problem. The content in Long CoT also demonstrates the formal reasoning abilities that emerge from our LoT-TL training pipeline in both prover and corrector agents. These observations qualitatively validate the multi-agent system’s design and emergence of formal reasoning ability in Long CoT, demonstrating its ability to combine high-level planning with iterative refinement in Lean-baed Long CoT.
@@ 4.1 Lean4 Theorem Proving using LLMs
The application of LLMs for FL proving has been a hot topic for study these years. The tree-search including works represented by Expert Iteration (Polu et al., 2022), ReProver (Yang et al., 2024b), Lean-Star (Lin et al., 2024), and InternLM-Step-Prover (Wu et al., 2024b). This direction does not take full consideration of the LLMs’ NL reasoning ability and costs exponentially increasing computation power. Another direction treats formal languages as code and asks the LLMs to do the complete-proof generation without interaction with the Lean executor to fully use the NL reasoning ability of LLMs. Significant works includes DeepSeek-Prover (Xin et al., 2024b, a), TheoremLlama (Wang et al., 2024), and Llemma (Azerbayev et al., 2023). Works in this direction tend to overlook the verification signals from Lean executors or do not have thorough thinking of the error messages.
4.2 Agent-based LLM
Traditional RL methods offer a training method solution for general reasoning and decision-making processes but often suffer from low sample efficiency and generalization problems (Pourchot & Sigaud, 2018). With the fast-developing reasoning and instruction-following ability of LLMs, many researchers began to make LLMs as agents (Wang et al., 2022; Ma et al., 2024). The primary method for using LLMs as agents is to design special prompts and in-context examples to let LLMs interact with the outsourcing tools using actionable responses (Xie et al., 2023; Yang et al., 2023). Further efforts were made to apply specialized training to enhance their agentic capabilities (Xu et al., 2023; Reed et al., 2022). In the context of formal reasoning, most tree-search methods (Yang et al., 2024b; Lin et al., 2024; Wu et al., 2024a) apply an LLM as an agent query the executor and receive feedback to further refine the proof. Because of the huge difference between FL and NL, such methods are unable to provide high-level analysis to the problem and provide a structured response.
@@ 5 Conclusion
This paper introduces MA-LoT, (to the best of our knowledge), the first multi-agent Lean-based Long Chain-of-Thought framework for formal theorem proving. Our approach addresses the limitations of single-agent systems, which either under-utilize the NL reasoning and planning capabilities of LLMs or fail to integrate formal verification feedback effectively. By structuring interactions between a prover and a corrector agent through Long CoT, MA-LoT enables deeper insights and long-term coherence in proof generation. To support this framework, we propose LoT-TL, a training pipeline that makes formal reasoning Long CoT capability emerge to LLMs without requiring annotated Lean Long CoT data. Through extensive experiments on the MiniF2F-Test benchmark, MA-LoT achieves 61.07% accuracy, surpassing all baselines, including both tree-search and whole-proof generation methods. These results underscore the advantages of integrating formal verification with structured reasoning leading to better Lean4 theorem proving capability. Beyond theorem proving, LoT-TL training pipeline demonstrates a potential method for applying Long CoT techniques to domain-specific tasks without specialized annotations. Additionally, the success of multi-agent Long CoT in Lean4 suggests broader applications of formal verification for enhancing structured reasoning across diverse fields.
@@ Appendix A Term chart
    MA-LoT (Multi-Agent Lean-based Long Chain-of-Thought framework): Our proposed multi-agent framework for formal theorem proving.
@@
    LoT-TL (LoT-Transfer Learning): The transfer learning pipeline we propose to enable LLMs with Lean4 Long CoT capability without the need for the specially annotated dataset.
@@
    LoT-Solver: The model we train through LoT-TL pipeline that serves both as prover and corrector agent. It is a high-standard Lean4 theorem proving model with Long CoT capability to control agent type and enhance formal thinking.
@@
    LoT-PD & LoT-CD (LoT-ProveData) & (LoT-CorrectionData): The LoT-PD is a dataset containing verified Lean4 theorem proofs and NL statement and proof. Its main function is to provide basic capability for the prover agent. The LoT-CD is a dataset containing incorrect error messages and correct Lean4 proof together with NL annotations. It majorly provides capability for the prover agent to refine proofs.
@@ Appendix C Scaling Law Study
Figure 3: Log-scale training steps and accuracy rate on MiniF2F of trained model. The dashed line is linear regression approximation line
@@
Appendix D Case studies
We present two case studies in Appendix D.1 and D.2. We can see that the multi-agent framework is able to write a more coherent and well-structured proof by making a high-level NL plan and executing such plan in the prover agent. The corrector agent is able to analyze the unsolved goals, make plans for correction, and write correct proof. This is achieved by our proposed training procedure.
D.1 Prover analysis
This is the example of provers agent generated result, in which we can see that the agent learns to write a high-level proof plan in the Long CoT and details execute the plan when writing the Lean4 code
@@ D.1 Prover analysis
D.2 Corrector analysis
This is an example of error correction that successfully solve an IMO problem, we can see that the model firstly analysis the error message, then make proof plan, and then write correct proof to another IMO-level problem
@@ D.2 Corrector analysis
Appendix E Training examples

### v3

@@ 3.6 Ablation Study
To show the effectiveness of each component of our LoT-TL training pipeline, we conduct this thorough ablation study. We demonstrate that the elements in LoT-TL pipeline work synergistically to strengthen the model’s formal theorem-proving capability through integrating FL in Long CoT. We use DeepSeek-Prover-v1.5-SFT as our base model and apply the pass@64 accuracy rate on the whole-proof generation method for this set of experiments. The results are presented in Table 3.
@@ 3.6.1 Effect of Training Stages
We evaluate training progression by measuring performance across key intermediate models, namely base model, base model + Long CoT, and base model + Long CoT + SFT, as shown in Table 3. Results indicate that training solely on NL CoT data provides minimal improvement. It suggests that the model lacks field-specific information to finish a comprehensive Lean4 proof under pure general NL Long CoT training. However, incorporating SFT data with the LoT-TL training method yields a marked improvement, demonstrating the effectiveness of transfer learning in equipping models with Lean4 Long CoT capabilities. Interestingly, although it is not designed for whole-proof writing, additional training with correction data further enhances the performance. This improvement likely arises from the model’s development of self-analysis capabilities in Lean4 code, allowing it to avoid potentially wrong solutions.
@@ 3.6.2 Switch-off Long CoT
This experiment shows that the strong FL reasoning power of LoT-Solver comes from the emergent formal reasoning ability in Long CoT, rather than trivially stacking more data. It uses our LoT-Solver model to write Lean4 proofs directly using the code completion method without Long CoT. We find the performance drop from 51.64% to 49.18% because the model does not take an explicit high-level plan in Long CoT, making it unable to finish some questions in the induction field.
@@ 3.6.3 Ablation of Long CoT
To validate the quality of integration between NL and FL in Long CoT, we fine-tune the DeepSeek-Prover-v1.5-RL model directly using our SFT dataset without Long CoT reasoning. We can find that the performance of w/o Long CoT model (48.36%) is lower than LoT-Solver (51.64%). The results confirm that Long CoT plays a crucial role in Lean4 theorem proving, offering structured reasoning that outperforms a direct RL-based additional data fine-tuned model.
@@ 3.7 Case Study
This section presents the general results of case studies of the MA-LoT framework. Due to the limited space, we leave detailed examples in Appendix E. From the examples, we can observe a clear collaboration between models. The prover can use a high-level plan to prove advanced MATH theorems, and the corrector can analyze the feedback from the Lean executor to formulate a correct proof of an IMO-level problem. The content in Long CoT also demonstrates the formal reasoning abilities that emerge from our LoT-TL training pipeline in both prover and corrector. These observations qualitatively validate the model-collaboration system’s design and emergence of formal reasoning ability in Long CoT, demonstrating its ability to combine high-level planning with iterative refinement in Lean-based Long CoT.
@@ 4.1 Lean4 Theorem Proving using LLMs
The application of LLMs for FL proving has been a hot topic for study in recent years. The tree-search including works represented by Expert Iteration (Polu et al., 2022), ReProver (Yang et al., 2024b), Lean-Star (Lin et al., 2024), and InternLM-Step-Prover (Wu et al., 2024b). This direction does not take full consideration of the LLMs’ NL reasoning ability and costs exponentially increasing computation power. Another direction treats formal languages as code and asks the LLMs to do the whole-proof generation without interaction with the Lean executor to fully use the NL reasoning ability of LLMs. Significant works includes DeepSeek-Prover (Xin et al., 2024b, a), TheoremLlama (Wang et al., 2024), and Llemma (Azerbayev et al., 2023). Works in this direction tend to overlook the verification signals from Lean executors or do not thoroughly think about the error messages. Although there is some early work (First et al., 2023; Zheng et al., 2023) that tries to combine both models, it did not perform well because of the lack of thorough NL reasoning and multi-round thinking in Long CoT.
4.2 LLM Collaboration
Traditional RL methods offer a training method solution for general reasoning and decision-making processes, but often suffer from low sample efficiency and generalization problems (Pourchot & Sigaud, 2018). With the fast-developing reasoning and instruction-following ability of LLMs, many researchers began to separate the cognition task for a single LLM  (Wang et al., 2022; Ma et al., 2024). The primary method for the separation of tasks is to design special prompts and in-context examples to let LLMs interact with the outsourcing tools using actionable responses (Xie et al., 2023; Yang et al., 2023). Further efforts were made to apply specialized training to enhance their field-specific capabilities (Xu et al., 2023; Reed et al., 2022). In the context of formal reasoning, most tree-search methods (Yang et al., 2024b; Lin et al., 2024; Wu et al., 2024a) apply an LLM to continuously query the executor and receive feedback to further refine the proof. Because of the huge difference between FL and NL, such methods are unable to provide a high-level analysis of the problem and provide a structured response.
@@ 5 Conclusion
This paper introduces MA-LoT, a comprehensive model-collaboration framework for formal theorem proving that leverages Lean-based Long Chain-of-Thought (CoT) reasoning. MA-LoT separate high-level proof planning from fine-grained correction, overcoming limitations of single-model approaches that either fail to harness LLMs’ NL reasoning capability or lack integration with formal feedback. By coordinating a prover and a corrector model through Long CoT, MA-LoT produces more structured, coherent, and insightful proofs. To support this framework, we introduce LoT-TL training-inference pipeline. This pipeline enables LLMs to develop domain-specific Long CoT reasoning capability through transfer learning without requiring specifically annotated Lean Long CoT data. We conduct extensive experiments on the MiniF2F benchmark, where the MA-LoT achieves 61.07% accuracy, surpassing all baselines, including both tree-search and whole-proof generation methods. These results highlight the value of combining formal verification with structured and iterative Long CoT reasoning. Beyond theorem proving, LoT-TL we proposed offers a general method for enabling Long CoT reasoning in other specialized domains without expensive RL or data annotation. Additionally, the success of the model-collaboration Long CoT in Lean4 suggests broader applications of formal verification for enhancing structured reasoning across diverse fields.
@@ Appendix A Term chart
    MA-LoT (Model-CollAboration Lean-based Long Chain-of-Thought framework): Our proposed model-collaboration framework for formal theorem proving.
@@
    LoT-TL (LoT-Transfer Learning): The transfer learning pipeline we propose to enable LLMs with Lean4 Long CoT capability without the need for a specially annotated dataset.
@@
    LoT-Solver: The model we train through LoT-TL pipeline that serves both as prover and corrector. It is a high-standard Lean4 theorem proving model with Long CoT capability to control the type of cognition task and enhance formal thinking.
@@
    LoT-PD & LoT-CD (LoT-ProveData) & (LoT-CorrectionData): The LoT-PD is a dataset containing verified Lean4 theorem proofs and NL statement and proof. Its main function is to provide basic capability for the prover. The LoT-CD is a dataset containing incorrect error messages and correct Lean4 proofs together with NL annotations. It mainly provides the capability for the prover model to refine proofs.
@@ Appendix C Scaling Law Study
Figure 3: Log-scale training steps and accuracy rate on MiniF2F of trained model. The dashed line is a linear regression approximation.
@@
Appendix D ProofNet result
We perform a small-scale baseline comparison on the ProofNet benchmark using the Goedel-Prover-SFT and MA-LoT(Goedel), where MA-LoT achieves 15.47% of accuracy rate under $16+16$ metric while the original Goedel-Prover is 12.15% under pass@32. The relatively low performance of both models is that Goedel-Prover does not fit the ProofNet well as reported in  Lin et al. (2025). However, we can still see the improvement of our method compared to the baseline, indicating MA-LoT’s effectiveness is not constrained to the baseline.
Appendix E Case studies
We present two case studies in Appendix E.1 and E.2. We can see that the model-collaboration framework can write a more coherent and well-structured proof by making a high-level NL plan and executing such a plan in the prover. The corrector can analyze the unsolved goals, make plans for correction, and write correct proof. This is achieved by our proposed training procedure.
E.1 Prover analysis
This is an example of a prover-generated result, in which we can see that the prover model learns to write a high-level proof plan in the Long CoT and details execute the plan when writing the Lean4 code
@@ E.1 Prover analysis
E.2 Corrector analysis
This is an example of error correction that successfully solves an IMO problem, We can see that the model first analyzes the error message, then makes a proof plan, and then writes a correct proof for another IMO-level problem
@@ E.2 Corrector analysis
Another example is demonstrated as follows:
@ Natural language theorem statement:mathd_algebra_419:What is the value of \$-a-b^2+3ab\$ if \$a=-1\$ and \$b=5\$? Show that it is -39.@ Lean4 theorem statement:```lean4theorem mathd_algebra_419  (a b : ℝ)  (h₀ : a = -1)  (h₁ : b = 5) :  -a - b^2 + 3 * (a * b) = -39 := by```&@ Lean4 theorem statement and proof with explanatory comments preceding each line of code:### Response:<Thought>Okay the user asks me to write Lean4 theorem proof using Long CoT, I should do the following:  1. Provide the natural language analysis for the theorem based on the Natural language theorem statement.  2. Draft the Lean4 tactics I should use to solve the problem  3. Write the output Lean4 code.The user also asks that I should avoid using the keyword `sorry` to give up the proof, so I will not write it in my Lean4 code.The natural language analysis for the theorem `mathd_algebra_419` is:Given a = -1 and b = 5, we need to compute the value of the expression -a - b² + 3ab and show that it equals -39.The steps to solve this problem are:1. Substitute a = -1 and b = 5 into the expression.2. Simplify the expression step by step.3. Verify that the result is -39.The Lean4 code for this theorem would involve substituting the given values for a and b into the expression and then simplifying it using algebraic operations.Now, let’s outline the Lean4 code:The theorem statement is:theorem mathd_algebra_419  (a b : ℝ)  (h₀ : a = -1)  (h₁ : b = 5) :  -a - b^2 + 3 * (a * b) = -39 := byThe proof would involve substituting the values of a and b into the expression and simplifying:First, substitute a = -1 and b = 5 into the expression:-a - b² + 3ab becomes:-(-1) - (5)² + 3*(-1)*(5)Simplify each term:-(-1) = 1-(5)² = -253*(-1)*(5) = -15So the expression becomes:1 - 25 - 15Now compute the result:1 - 25 = -24-24 -15 = -39Thus, the value of the expression is -39.Now, let’s write the Lean4 code.The code would start with the theorem statement, then use the `rw` tactic to substitute the values of a and b.Wait, but in Lean4, we can directly substitute the values into the expression.So the code would be:theorem mathd_algebra_419  (a b : ℝ)  (h₀ : a = -1)  (h₁ : b = 5) :  -a - b^2 + 3 * (a * b) = -39 := by  rw [h₀, h₁]  ringBut wait, let me check. The expression is -a - b² + 3ab. So substituting a = -1 and b = 5:-(-1) - 5² + 3*(-1)*5 = 1 -25 -15 = -39.Yes, that’s correct.So the code can be written as:rw [h₀, h₁]Then use the `ring` tactic to simplify the expression.So the complete code would be:theorem mathd_algebra_419  (a b : ℝ)  (h₀ : a = -1)  (h₁ : b = 5) :  -a - b^2 + 3 * (a * b) = -39 := by  rw [h₀, h₁]  ringAlternatively, using the `simp` tactic with the appropriate lemmas.But in this case, the `ring` tactic is sufficient.So the code is straightforward.I should test this code in Lean4 to ensure that it works.But since the user asked to avoid using `sorry`, I will not write the code. Instead, I will provide the explanation.</Thought><Output>To solve the problem, we need to evaluate the expression \(-a - b^2 + 3ab\) given \(a = -1\) and \(b = 5\), and show that the result is \(-39\).### Key steps:1. **Substitute the given values into the expression:**   - Replace \(a\) with \(-1\) and \(b\) with \(5\).   - The expression becomes \(-(-1) - (5)^2 + 3(-1)(5)\).2. **Simplify each term:**   - \(-(-1) = 1\)   - \((5)^2 = 25\)   - \(3(-1)(5) = -15\)3. **Combine the simplified terms:**   - \(1 - 25 - 15 = -39\)### Proof:```lean4theorem mathd_algebra_419  (a b : ℝ)  (h₀ : a = -1)  (h₁ : b = 5) :  -a - b^2 + 3 * (a * b) = -39 := by  rw [h₀, h₁]  ring```&### Explanation:- **Substitution:** The `rw [h₀, h₁]` tactic replaces \(a\) with \(-1\) and \(b\) with \(5\) in the expression.- **Simplification:** The `ring` tactic simplifies the expression using algebraic rules, resulting in \(-39\).This approach ensures that we efficiently and correctly evaluate the expression by substituting the given values and simplifying step-by-step. The final result confirms that the value of the expression is indeed \(-39\).</Output>
From the example above, we can clearly see the saying of wait, which indicates a rethinking process of the Long CoT. Such rethinking can help the model to reach conclusion that needs more thorough thinking.
E.3 Theorem only solved by MA-LoT
We have observed that algebra_2varlineareq_fp3zeq11_3tfm1m5zeqn68_feqn10_zeq7 is only solved in the MA-LoT setup. From the original code draft and the error analysis, we can spot that our corrector framework is able to restructure the proof from the previous left and right cases into direct solving. This indicates our method’s superiority in rethinking based on formal guidance.
<Thought>Alright, I need to prove the theorem lean_workbook_plus_68493 using the Lean4 code. Here is my draft of the proof:```lean4import Mathlibimport Aesopset_option maxHeartbeats 0open BigOperators Real Nat Topology Rattheorem algebra_2varlineareq_fp3zeq11_3tfm1m5zeqn68_feqn10_zeq7  (f z: ℂ)  (h₀ : f + 3*z = 11)  (h₁ : 3*(f - 1) - 5*z = -68) :  f = -10 ∧ z = 7 := by  /-  Given the equations \( f + 3z = 11 \) and \( 3(f - 1) - 5z = -68 \), we need to show that \( f = -10 \) and \( z = 7 \).  1. Start by isolating \( f \) from the first equation: \( f = 11 - 3z \).  2. Substitute \( f = 11 - 3z \) into the second equation: \( 3((11 - 3z) - 1) - 5z = -68 \).  3. Simplify the equation step by step:     - Distribute the 3: \( 3(10 - 3z) - 5z = -68 \).     - Distribute the 3 again: \( 30 - 9z - 5z = -68 \).     - Combine like terms: \( 30 - 14z = -68 \).  4. Solve for \( z \):     - Subtract 30 from both sides: \( -14z = -98 \).     - Divide both sides by -14: \( z = 7 \).  5. Substitute \( z = 7 \) back into \( f = 11 - 3z \):     - \( f = 11 - 3(7) \).     - \( f = 11 - 21 \).     - \( f = -10 \).  Thus, we have shown that \( f = -10 \) and \( z = 7 \).  -/  constructor  case left =>    -- Isolate f from the first equation    have h₂ : f = 11 - 3*z := by linear_combination h₀    -- Substitute f into the second equation    have h₃ : 3*(11 - 3*z - 1) - 5*z = -68 := by rw [h₂] at h₁; exact h₁    -- Simplify the equation to solve for z    have h₄ : 30 - 14*z = -68 := by linear_combination h₃    have h₅ : -14*z = -98 := by linear_combination h₄    have h₆ : z = 7 := by linear_combination h₅ / -14    exact h₆  case right =>    -- Isolate z from the first equation    have h₂ : z = 7 := by linear_combination h₀    exact h₂```&Let me test it in Lean4Emmm, it seems the above proof is wrong.Let me check the error messages.OK, Here is the error messages:```bashline 39type mismatch  h₆has type  z = 7 : Propbut is expected to have type  f = -10 : Prop```&```bashline 42ring failed, ring expressions not equalf z : ℂh₀ : f + 3 * z = 11h₁ : 3 * (f - 1) - 5 * z = -68⊢ 4 + (-(z * 2) - f) = 0```&So, I will rethink a Lean4 proof following the steps  1. Provide the natural language analysis for the theorem based on the Natural language theorem statement, Lean4 theorem statement, my previous proof and the error message.  2. Draft the Lean4 tactics I should use to solve the problem  3. Write the output Lean4 code.Let me analysis the wrong Lean4 solution through the error messagesThe first error says:line 39: type mismatchh₆ has type z = 7 : Prop but is expected to have type f = -10 : PropWait, in the code, I have two cases: left and right. The left case is supposed to prove f = -10, and the right case z = 7. But in the left case, I derived z = 7 and then tried to use h₆ to conclude f = -10, but h₆ is z = 7. That’s the type mismatch. So, I need to fix that.The second error is:line 42: ring failed, ring expressions not equalf z : ℂh₀ : f + 3 * z = 11h₁ : 3 * (f - 1) - 5 * z = -68⊢ 4 + (-(z * 2) - f) = 0Hmm, this seems like an unrelated expression that Lean is trying to check. Maybe there’s a mistake in how I’m handling the substitutions or the algebraic manipulations.Let me revisit the code. In the left case (proving f = -10), I first derived f = 11 - 3*z from h₀, then substituted that into h₁ to get h₃. Then I simplified h₃ to get h₄, h₅, and h₆, which led to z = 7. But then I tried to use h₆ to conclude f = -10, but h₆ is z = 7. I need to instead use z = 7 to find f.So, in the left case, after getting z = 7, I should substitute that back into f = 11 - 3*z to get f = -10. Then I can use that to conclude f = -10.Similarly, in the right case, I need to properly handle the substitution and simplification steps to avoid the ring error.Let me try adjusting the code accordingly. Instead of having separate cases for left and right, maybe I should first solve for z, then use that to solve for f, and then combine the results. Let’s see how the corrected code does it.</Thought><Output>```lean4import Mathlibimport Aesopset_option maxHeartbeats 0open BigOperators Real Nat Topology Rat/--Given that \$f + 3z = 11\$ and \$3(f - 1) - 5z = -68\$, show that \$f = -10\$ and \$z = 7\$.-/theorem algebra_2varlineareq_fp3zeq11_3tfm1m5zeqn68_feqn10_zeq7  (f z: ℂ)  (h₀ : f + 3*z = 11)  (h₁ : 3*(f - 1) - 5*z = -68) :  f = -10 ∧ z = 7 := by  /-  Given the equations \( f + 3z = 11 \) and \( 3(f - 1) - 5z = -68 \), we aim to show that \( f = -10 \) and \( z = 7 \).  1. Start by isolating \( f \) in the first equation: \( f = 11 - 3z \).  2. Substitute \( f = 11 - 3z \) into the second equation: \( 3(11 - 3z - 1) - 5z = -68 \).  3. Simplify the equation step by step:     - Expand the terms inside the parentheses: \( 3(10 - 3z) - 5z = -68 \).     - Distribute the 3: \( 30 - 9z - 5z = -68 \).     - Combine like terms: \( 30 - 14z = -68 \).  4. Solve for \( z \):     - Subtract 30 from both sides: \( -14z = -98 \).     - Divide both sides by -14: \( z = 7 \).  5. Substitute \( z = 7 \) back into \( f = 11 - 3z \):     - Calculate \( f \): \( f = 11 - 3(7) = 11 - 21 = -10 \).  6. Conclude that \( f = -10 \) and \( z = 7 \).  -/  -- Isolate f in the first equation  have h₂ : f = 11 - 3*z := by linear_combination h₀  -- Substitute f = 11 - 3z into the second equation  have h₃ : 3*(11 - 3*z - 1) - 5*z = -68 := by rw [h₂] at h₁; exact h₁  -- Simplify the equation step by step  have h₄ : 3*10 - 9*z - 5*z = -68 := by linear_combination h₃  have h₅ : 30 - 14*z = -68 := by linear_combination h₄  -- Solve for z  have h₆ : -14*z = -98 := by linear_combination h₅ - 30  have h₇ : z = 7 := by linear_combination h₆ / -14  -- Substitute z = 7 back into f = 11 - 3z  have h₈ : f = -10 := by rw [h₇] at h₂; linear_combination h₂  -- Conclude the proof  exact ⟨h₈, h₇⟩```&</Output>
Appendix F Training examples